# Higher property T and below-rank phenomena of lattices

### v1

@@ 2.5. Hilbert $C^{\ast}$-modules
Let $A$ be a $C^{\ast}$-algebra. A Hilbert $A$-module is an $A$-module $E$ with an $A$-valued inner product $\langle\_,\_\rangle\colon E\times E\to A$ such that $E$ is complete with respect to the ($\mathbb{C}$}-valued) norm ${\left\lVert x\right\rVert}={\left\lVert\langle x,x\rangle\right\rVert}_{A}^{1/2}$. We refer to the book [lance] as a background reference.
The category of Hilbert $A$-modules with bounded $A$-linear maps as morphisms for an arbitrary $C^{\ast}$-algebra $A$ is very similar to the special case $A={\mathbb{C}}$ which is just about Hilbert spaces and bounded operators. But there are some important differences. A closed $A$-submodule $F\subset E$ is complemented if $E=F\oplus F^{\perp}$ where $F^{\perp}$ is defined in the obvious sense. Not every closed $A$-submodule is complemented. Further, not every bounded $A$-linear map has an adjoint. We usually work in the category of Hilbert $A$-modules with bounded and adjointable $A$-linear maps as morphisms. If $V,W$ are Hilbert $A$-modules, we denote the set of bounded adjointable maps from $V$ to $W$ by $B(V,W)$. Furthermore, $B(V,V)$ is a $C^{\ast}$-algebra [lance]*p. 8. So a bounded adjointable operator $f\colon V\to V$ is positive if it is positive as an element of $B(V,V)$. This is implied by the positivity $\langle f(x),x\rangle\geq 0$ of the $A$-valued inner product by [lance]*Lemma 4.1.
Let $E$ be a finitely generated projective $A$-module in the algebraic sense [^9] In the case of a commutative $C^{\ast}$-algebra $C(X)$ such an $E$ corresponds to a finite-dimensional vector bundle over $X$.. Then $E$ embeds as a direct summand of some finitely generated free module $A^{n}$ and inherits the structure of a Hilbert $A$-module from that embedding. The Hilbert $A$-structure is independent of the choice up to unitary isomorphism. This follows from combining [lance]*Corollary 3.3 and Proposition 3.8. Furthermore, every $A$-linear map between finitely generated projective $A$-modules is bounded and adjointable. In case of finitely generated free $A$-modules the $A$-linear map $f$ is represented by a matrix $M$, and the adjoint of $f$ is represented by the matrix obtained from $M$ by taking the transpose and the element-wise adjoint. In the general case one uses [lance]*Corollary 3.3 again.
@@
Ad $(3)\Rightarrow(2)$. We refer to the setting in §2.2 but drop the superscript $A$ in the notation of the differentials. The vanishing of $H_{n}(\Gamma,A)$ follows from Lemma 21. By [lance]*Proposition 3.2 and its proof the subspaces $\ker(d_{n})$ and $\ker(d_{n+1}^{\ast})$ of $P_{n}^{A}$ are complementable. Further, $\ker(d_{n})^{\perp}$ is the closure of $\operatorname{im}(d_{n}^{\ast})$, and $\ker(d_{n+1}^{\ast})^{\perp}$ is the closure of $\operatorname{im}(d_{n+1})$. Using that one shows the orthogonal decomposition
@@
Ad $(2)\Rightarrow(3)$. First we show that $\Delta_{n}$ is injective. Since $\operatorname{im}(d_{n+1})=\ker(d_{n})$ is closed, $\operatorname{im}(d_{n+1})$ is complementable by [lance]*Proposition 3.2. Further, $\ker(d_{n+1}^{\ast})$ is the orthogonal complement of $\operatorname{im}(d_{n+1})$ [lance]*p. 23. Therefore $\ker(d_{n})\cap\ker(d_{n+1}^{\ast})$ injects into $H_{(}\Gamma,A)$. Thus $\ker(d_{n})\cap\ker(d_{n+1}^{\ast})=0$. On the other hand, we have $\langle\Delta_{n}x,x\rangle={\left\lVert d_{n}x\right\rVert}^{2}+{\left\lVert d_{n+1}^{\ast}x\right\rVert}^{2}$. This implies that $\Delta_{n}$ is a positive operator [lance]*Lemma 4.1 from which we can conclude that $\Delta_{n}$ has a square root. Hence $\langle\Delta_{n}x,x\rangle=0$ implies that $x=0$ and we obtain that
@@ 2.6. Unitary $\Gamma$-modules
The first statement is well known. See, for example, [Lueck, Lemma 1.18] or [Bader-Nowak, Proposition 16.1]. The isomorphism $\bar{H}_{n}(\Gamma,V)\cong\ker\Delta_{n}^{V}$ follows from the orthogonal decompositions
@@
Similarly for the cohomological version. The equivalence (1) $\Leftrightarrow$ (2) follows by Lemma 23 and the fact that $V\cong V^{\ast}$ as ${\mathbb{C}}\Gamma$-modules. Note that if $V$ is a right ${\mathbb{C}}\Gamma$-module, then $V^{\ast}$ is a left ${\mathbb{C}}\Gamma$-module and vice versa. So in the isomorphism we implicitly use the involution on ${\mathbb{C}}\Gamma$ to regard $V$ as a left ${\mathbb{C}}\Gamma$-module. Lemma 21(3) gives (3) $\Rightarrow$ (1) and (2), see Remark 20. On the other hand (2) implies that both maps $d_{n}$ and $d_{n+1}$ have closed ranges, so also their adjoints $d_{n}^{*}$ and $d_{n+1}^{*}$ have closed ranges, and so do the compositions $d_{n}^{*}d_{n}$ and $d_{n+1}d_{n+1}^{*}$. Since the images of these maps are orthogonal, we get that also $\Delta_{n}$ has a closed image. Since $\Delta_{n}$ is a positive operator injective operator, we conclude that it is invertible, thus (2) $\Rightarrow$ (3). Similarly, (1) $\Rightarrow$ (3). See also see [Bader-Nowak, Proposition 16.5] for (1) $\Leftrightarrow$ (3). ∎
@@ 2.7. Universal envelopes
The forgetful functor from the category of von Neumann algebras to the category of $C^{\ast}$-algebras has a left adjoint called the von Neumann envelope functor. By a result of Sherman-Takeda the von Neumann envelope of a $C^{\ast}$-algebra is its double dual (as a von Neumann algebra). See [takesaki]*III.2.
@@ 2.8. The augmentation ideals and the Kazhdan projection
Lemma 38 ([Akemann-Walter]*Lemma 2).
$\Gamma$ has property T if and only if the Kazhdan projection $z$ is in $C^{*}\Gamma$.
@@
If $\Gamma$ has property T, then the trivial representation is isolated in its unitary dual. Hence it corresponds to a central projection in $C^{*}\Gamma$, which must be $z$, the unique invariant central projection in $W^{*}\Gamma$. If $z\in C^{*}\Gamma$, then the class of unitary representation of $\Gamma$ that have no non-trivial invariants coincides with the class of *-representation of the unital algebra $C_{0}^{*}\Gamma$, which is closed under taking ultrapowers. By Lemma 33 and Lemma 30 such representations have no almost invariant vectors. It follows that $\Gamma$ has property T. ∎
@@
By Lemma 24, if $\Gamma$ does not have property T then $H_{0}$ and $H^{1}$ are non-Hausdorff also for $W^{*}\Gamma$ coefficients. Using the long exact sequences associated with the maps $\epsilon$, one sees that this is also the case for $C_{0}^{*}\Gamma$ and $W_{0}^{*}\Gamma$ coefficients.
@@
Using Lemma 23 we will show that $H_{0}(\Gamma,C^{*}\Gamma)$ is not Hausdorff by showing that $H^{1}(\Gamma,(C^{*}\Gamma)^{*})$ is not Hausdorff. The restriction to $\Gamma\subset C^{*}\Gamma$ gives an embedding $(C^{*}\Gamma)^{*}\hookrightarrow\ell^{\infty}(\Gamma)$ whose image is the Fourier-Stieltjes algebra of all matrix coefficients. By Eymard’s theorem [eymard], the state space of $C^{*}\Gamma$ is mapped onto the space $\text{PD}_{1}(\Gamma)$ of positive definite functions of norm $1$ under the embedding. Further, the restriction of the weak*-topology from both spaces coincide on $\text{PD}_{1}(\Gamma)$.
@@ 2.9. C*-algebraic reformulation of higher T
For $T_{n}]$ the theorem follows at once from Lemma 35 upon setting $A=C^{*}\Gamma$. We now consider the proof for $(T_{n})$. Assume first that $\Gamma$ does not have $T$. Then, for every $n>0$, none of the conditions (1)-(3) is satisfied, by Lemma 40 and Remark 41. We thus assume as we may that $\Gamma$ has T, and in view of Lemma 38, we apply Convention 39, thus considering $C_{0}^{*}\Gamma$ as a unital C*-algebra extending ${\mathbb{C}}\Gamma$, such that every unitary representation of $\Gamma$ which has no non-trivial invariants extends to a *-representation of $C_{0}^{*}\Gamma$ and vice versa. The theorem now follows again from Lemma 35 upon setting $A=C_{0}^{*}\Gamma$.
@@
The only part which does not follow immediately from Lemma 35 is the last statement. However, using the fact that $H^{n+1}(\Gamma,{\mathbb{C}})$ is finite dimensional by Convention 17, we get
@@
We now turn to consider von Neumann algebras. Given a group $\Gamma$, a von Neumann algebra $M$ and a group homomorphism $\Gamma\to\mathcal{U}(M)$ to the group of unitary elements of $M$, we consider $M$ as a ${\mathbb{C}}\Gamma$-module with respect to the left action. We call such modules $\Gamma$-unitary von Neumann algebras. Every von Neumann algebra $M$ has a unique predual, which we denote $M_{*}$, and a one-parameter collection of $M$-bimodules, interpolating between $M_{*}$ and $M$ which we now consider. For $1\leq p\leq\infty$, there is an associated non-commutative $L^{p}$-space $L^{p}(M)$, where $L^{\infty}(M)\cong M$ and $L^{1}(M)\cong M_{*}$. For $1<p<\infty$, $L^{p}(M)$ is super-reflexive and $L^{p}(M)^{*}\cong L^{q}(M)$ where $1/p+1/q=1$. In particular, $L^{2}(M)$ is a Hilbert space. For more information, see[Pisier-Xu].
@@
We note that the group homomorphism $\Gamma\to\mathcal{U}(M)$ extends to a C*-algebra morphism $C^{*}\Gamma\to M$, thus every $M$-module is a $C^{*}\Gamma$-module. The first part, regarding $T_{n}]$ groups, follows from Theorem 12. For the second part, regarding $(T_{n})$ groups, use also Lemma 42. Note that if the map $C^{*}\Gamma\to M$ factors via $C_{0}^{*}\Gamma$ then all the relevant (co)homologies vanish. In particular, for every $p$, $L^{p}(M)^{\Gamma}=H^{0}(\Gamma,L^{p}(M))=0$. By Lemma 42, this is the case if $M^{\Gamma}=0$. Assume now that for some $1\leq p<\infty$, $L^{p}(M)^{\Gamma}=0$. By considering the Mazur map as in [Olivier], we assume as we may $p=1$, that is $(M_{*})^{\Gamma}=0$. We conclude that $M_{*}$ is a $C_{0}^{*}\Gamma$-module, and by duality $M$ is a $C_{0}^{*}\Gamma$-module. This concludes the proof. ∎
@@
This conjecture in degree 1 is due to Mimura. Then it is known to hold for $M=B(H)$ and $1<p\leq 2$ by [Mikala, Theorem 3(1)] and for any $M$ commutative by [BFGM] ($1<p\leq 2$) and [BGM] ($p=1$). If $p>2$, the analogous conjecture is known not to hold, already in degree 1 and for $M$ commutative on which $\Gamma$ acts by automorphisms. However, we expect it to hold assuming the existence of an invariant state.
@@
In degree 1 and for $2\leq p<\infty$, this conjecture is proved by Marrakchi and de la Salle, [Mikala, Theorem 3.(2)]. Note that for the profinite completion $\hat{\Gamma}$, we have $H^{1}(\Gamma,L^{\infty}(\hat{\Gamma}))\neq 0$, so the conjecture could not be extended to $p=\infty$ [^10] We thank Mikael de la Salle for sharing this observation with us..

### v2

@@ 2.5. Hilbert $C^{\ast}$-modules
Let $A$ be a $C^{\ast}$-algebra. A Hilbert $A$-module is an $A$-module $E$ with an $A$-valued inner product $\langle\_,\_\rangle\colon E\times E\to A$ such that $E$ is complete with respect to the ($\mathbb{C}$}-valued) norm ${\left\lVert x\right\rVert}={\left\lVert\langle x,x\rangle\right\rVert}_{A}^{1/2}$. We refer to the book [67] as a background reference.
The category of Hilbert $A$-modules with bounded $A$-linear maps as morphisms for an arbitrary $C^{\ast}$-algebra $A$ is very similar to the special case $A={\mathbb{C}}$ which is just about Hilbert spaces and bounded operators. But there are some important differences. A closed $A$-submodule $F\subset E$ is complemented if $E=F\oplus F^{\perp}$ where $F^{\perp}$ is defined in the obvious sense. Not every closed $A$-submodule is complemented. Further, not every bounded $A$-linear map has an adjoint. We usually work in the category of Hilbert $A$-modules with bounded and adjointable $A$-linear maps as morphisms. If $V,W$ are Hilbert $A$-modules, we denote the set of bounded adjointable maps from $V$ to $W$ by $B(V,W)$. Furthermore, $B(V,V)$ is a $C^{\ast}$-algebra [67]*p. 8. So a bounded adjointable operator $f\colon V\to V$ is positive if it is positive as an element of $B(V,V)$. This is implied by the positivity $\langle f(x),x\rangle\geq 0$ of the $A$-valued inner product by [67]*Lemma 4.1.
Let $E$ be a finitely generated projective $A$-module in the algebraic sense [^9] In the case of a commutative $C^{\ast}$-algebra $C(X)$ such an $E$ corresponds to a finite-dimensional vector bundle over $X$.. Then $E$ embeds as a direct summand of some finitely generated free module $A^{n}$ and inherits the structure of a Hilbert $A$-module from that embedding. The Hilbert $A$-structure is independent of the choice up to unitary isomorphism. This follows from combining [67]*Corollary 3.3 and Proposition 3.8. Furthermore, every $A$-linear map between finitely generated projective $A$-modules is bounded and adjointable. In case of finitely generated free $A$-modules the $A$-linear map $f$ is represented by a matrix $M$, and the adjoint of $f$ is represented by the matrix obtained from $M$ by taking the transpose and the element-wise adjoint. In the general case one uses [67]*Corollary 3.3 again.
@@
Ad $(3)\Rightarrow(2)$. We refer to the setting in §2.2 but drop the superscript $A$ in the notation of the differentials. The vanishing of $H_{n}(\Gamma,A)$ follows from Lemma 26. By [67]*Proposition 3.2 and its proof the subspaces $\ker(d_{n})$ and $\ker(d_{n+1}^{\ast})$ of $P_{n}^{A}$ are complementable. Further, $\ker(d_{n})^{\perp}$ is the closure of $\operatorname{im}(d_{n}^{\ast})$, and $\ker(d_{n+1}^{\ast})^{\perp}$ is the closure of $\operatorname{im}(d_{n+1})$. Using that one shows the orthogonal decomposition
@@
Ad $(2)\Rightarrow(3)$. First we show that $\Delta_{n}$ is injective. Since $\operatorname{im}(d_{n+1})=\ker(d_{n})$ is closed, $\operatorname{im}(d_{n+1})$ is complementable by [67]*Proposition 3.2. Further, $\ker(d_{n+1}^{\ast})$ is the orthogonal complement of $\operatorname{im}(d_{n+1})$ [67]*p. 23. Therefore $\ker(d_{n})\cap\ker(d_{n+1}^{\ast})$ injects into $H_{(}\Gamma,A)$. Thus $\ker(d_{n})\cap\ker(d_{n+1}^{\ast})=0$. On the other hand, we have $\langle\Delta_{n}x,x\rangle={\left\lVert d_{n}x\right\rVert}^{2}+{\left\lVert d_{n+1}^{\ast}x\right\rVert}^{2}$. This implies that $\Delta_{n}$ is a positive operator [67]*Lemma 4.1 from which we can conclude that $\Delta_{n}$ has a square root. Hence $\langle\Delta_{n}x,x\rangle=0$ implies that $x=0$ and we obtain that
@@ 2.6. Unitary $\Gamma$-modules
The first statement is well known. See, for example, [75, Lemma 1.18] or [13, Proposition 16.1]. The isomorphism $\bar{H}_{n}(\Gamma,V)\cong\ker\Delta_{n}^{V}$ follows from the orthogonal decompositions
@@
Similarly for the cohomological version. The equivalence (1) $\Leftrightarrow$ (2) follows by Lemma 28 and the fact that $V\cong V^{\ast}$ as ${\mathbb{C}}\Gamma$-modules. Note that if $V$ is a right ${\mathbb{C}}\Gamma$-module, then $V^{\ast}$ is a left ${\mathbb{C}}\Gamma$-module and vice versa. So in the isomorphism we implicitly use the involution on ${\mathbb{C}}\Gamma$ to regard $V$ as a left ${\mathbb{C}}\Gamma$-module. Lemma 26(3) gives (3) $\Rightarrow$ (1) and (2), see Remark 25. On the other hand (2) implies that both maps $d_{n}$ and $d_{n+1}$ have closed ranges, so also their adjoints $d_{n}^{*}$ and $d_{n+1}^{*}$ have closed ranges, and so do the compositions $d_{n}^{*}d_{n}$ and $d_{n+1}d_{n+1}^{*}$. Since the images of these maps are orthogonal, we get that also $\Delta_{n}$ has a closed image. Since $\Delta_{n}$ is a positive operator injective operator, we conclude that it is invertible, thus (2) $\Rightarrow$ (3). Similarly, (1) $\Rightarrow$ (3). See also see [13, Proposition 16.5] for (1) $\Leftrightarrow$ (3). ∎
@@ 2.7. Universal envelopes
The forgetful functor from the category of von Neumann algebras to the category of $C^{\ast}$-algebras has a left adjoint called the von Neumann envelope functor. By a result of Sherman-Takeda the von Neumann envelope of a $C^{\ast}$-algebra is its double dual (as a von Neumann algebra). See [95]*III.2.
@@ 2.8. The augmentation ideals and the Kazhdan projection
Lemma 43 ([5]*Lemma 2).
A group $\Gamma$ has property T if and only if the Kazhdan projection $z$ is in $C^{*}\Gamma$.
@@
If $\Gamma$ has property T, then the trivial representation is isolated in its unitary dual. Hence it corresponds to a central projection in $C^{*}\Gamma$, which must be $z$, the unique invariant central projection in $W^{*}\Gamma$. If $z\in C^{*}\Gamma$, then the class of unitary representations of $\Gamma$ that have no non-trivial invariants coincides with the class of *-representation of the unital algebra $C_{0}^{*}\Gamma$, which is closed under taking ultrapowers. By Lemma 38 and Lemma 35 such representations have no almost invariant vectors. It follows that $\Gamma$ has property T. ∎
@@
If $\Gamma$ does not have property T then Lemma 28 implies that $H_{0}(\Gamma,W^{*}\Gamma)$ is non-Hausdorff and using Lemma 29 we also get that $H^{1}(\Gamma,W^{*}\Gamma)$ is non-Hausdorff. Using the long exact sequences associated with the maps $\epsilon$, one sees that $H_{0}$ and $H^{1}$ are non-Hausdorff also for $C_{0}^{*}\Gamma$ and $W_{0}^{*}\Gamma$ coefficients.
@@
Using Lemma 28 we will show that $H_{0}(\Gamma,C^{*}\Gamma)$ is not Hausdorff by showing that $H^{1}(\Gamma,(C^{*}\Gamma)^{*})$ is not Hausdorff. The restriction to $\Gamma\subset C^{*}\Gamma$ gives an embedding $(C^{*}\Gamma)^{*}\hookrightarrow\ell^{\infty}(\Gamma)$ whose image is the Fourier-Stieltjes algebra of all matrix coefficients. By Eymard’s theorem [42], the state space of $C^{*}\Gamma$ is mapped onto the space $\text{PD}_{1}(\Gamma)$ of positive definite functions of norm $1$ under the embedding. Further, the restriction of the weak*-topology from both spaces coincide on $\text{PD}_{1}(\Gamma)$.
@@
Remark 47.
The injection $(C^{*}\Gamma)^{*}\hookrightarrow\ell^{\infty}(\Gamma)$ alluded to above shows that $((C^{*}\Gamma)^{*})^{\Gamma}$ is one dimensional, spanned by $\epsilon$.
@@ 2.9. C*-algebraic reformulation of higher T
For $T_{n}]$ the theorem follows at once from Lemma 40 upon setting $A=C^{*}\Gamma$. We now consider the proof for $(T_{n})$. Assume first that $\Gamma$ does not have $T$. Then, for every $n>0$, none of the conditions (1)-(3) is satisfied, by Lemma 45 and Remark 46. We thus assume as we may that $\Gamma$ has T, and in view of Lemma 43, we apply Convention 44, thus considering $C_{0}^{*}\Gamma$ as a unital C*-algebra extending ${\mathbb{C}}\Gamma$, such that every unitary representation of $\Gamma$ which has no non-trivial invariants extends to a *-representation of $C_{0}^{*}\Gamma$ and vice versa. The theorem now follows again from Lemma 40 upon setting $A=C_{0}^{*}\Gamma$.
@@
The only part which does not follow immediately from Lemma 40 is the last statement. However, using the fact that $H^{n+1}(\Gamma,{\mathbb{C}})$ is finite dimensional by Convention 22, we get
@@
We now turn to consider von Neumann algebras. Given a group $\Gamma$, a von Neumann algebra $M$ and a group homomorphism $\Gamma\to\mathcal{U}(M)$ to the group of unitary elements of $M$, we consider $M$ as a ${\mathbb{C}}\Gamma$-module with respect to the left action. We call such modules $\Gamma$-unitary von Neumann algebras. Every von Neumann algebra $M$ has a unique predual, which we denote $M_{*}$, and a one-parameter collection of $M$-bimodules, interpolating between $M_{*}$ and $M$ which we now consider. For $1\leq p\leq\infty$, there is an associated non-commutative $L^{p}$-space $L^{p}(M)$, where $L^{\infty}(M)\cong M$ and $L^{1}(M)\cong M_{*}$. For $1<p<\infty$, $L^{p}(M)$ is super-reflexive and $L^{p}(M)^{*}\cong L^{q}(M)$ where $1/p+1/q=1$. In particular, $L^{2}(M)$ is a Hilbert space. For more information, see[88].
@@
We note that the group homomorphism $\Gamma\to\mathcal{U}(M)$ extends to a C*-algebra morphism $C^{*}\Gamma\to M$, thus every $M$-module is a $C^{*}\Gamma$-module. The first part, regarding $T_{n}]$ groups, follows from Theorem 17. For the second part, regarding $(T_{n})$ groups, use also Lemma 48. Note that if the map $C^{*}\Gamma\to M$ factors via $C_{0}^{*}\Gamma$ then all the relevant (co)homologies vanish. In particular, for every $p$, $L^{p}(M)^{\Gamma}=H^{0}(\Gamma,L^{p}(M))=0$. By Lemma 48, this is the case if $M^{\Gamma}=0$. Assume now that for some $1\leq p<\infty$, $L^{p}(M)^{\Gamma}=0$. By considering the Mazur map as in [80], we assume as we may $p=1$, that is $(M_{*})^{\Gamma}=0$. We conclude that $M_{*}$ is a $C_{0}^{*}\Gamma$-module, and by duality $M$ is a $C_{0}^{*}\Gamma$-module. This concludes the proof. ∎
@@
This conjecture in degree 1 is due to Mimura. Then it is known to hold for $M=B(H)$ and $1<p\leq 2$ by [76, Theorem 3(1)] and for any $M$ commutative by [10] ($1<p\leq 2$) and [8] ($p=1$). If $p>2$, the analogous conjecture is known not to hold, already in degree 1 and for $M$ commutative on which $\Gamma$ acts by automorphisms. However, we expect it to hold assuming the existence of an invariant state.
@@
In degree 1 and for $2\leq p<\infty$, this conjecture is proved by Marrakchi and de la Salle, [76, Theorem 3.(2)]. Note that for the profinite completion $\hat{\Gamma}$, we have $H^{1}(\Gamma,L^{\infty}(\hat{\Gamma}))\neq 0$, so the conjecture could not be extended to $p=\infty$ [^10] We thank Mikael de la Salle for sharing this observation with us..